OpenStudy (anonymous):

Find the slope of the line that contains these points. (3, 2) (5, 12)

OpenStudy (anonymous):

Given (3, 2) = (x1, y1) and (5, 12)= (x2, y2) slope of line is given as: m = (y2-y1)/(x2-x1) =(12-2)/(5-3)=(10)/(2) = 5